OpenStudy (anonymous):

How many would I go over and up If I was graphing 2x-3y=12

ganeshie8 (ganeshie8):

rearrange it in \(y = mx + b\) form

OpenStudy (anonymous):

Can you please show me how to do that please?

ganeshie8 (ganeshie8):

\(2x-12 = 3y \) \(y = (2/3)x - 4\)

ganeshie8 (ganeshie8):

slope = rise/run = 2/3

ganeshie8 (ganeshie8):

so you need to RISE 2 units, after running for 3 units right

OpenStudy (anonymous):

Ok thanks alot!.

ganeshie8 (ganeshie8):

you wlc :)